40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 1 08:50:48 2022, yuta, Summary, LSC, FPMI with REFL/AS55 trial 
    Reply  Thu Jul 14 00:10:08 2022, yuta, Summary, LSC, FPMI with REFL/AS55 trial continued AS55_I.pngAS55_Q.pngREFL55_I.pngREFL55_Q.png
       Reply  Fri Jul 15 19:13:22 2022, Paco, Summary, LSC, FPMI with REFL/AS55 demod phase adjust 
          Reply  Fri Jul 15 22:36:04 2022, rana, Summary, LSC, FPMI with REFL/AS55 demod phase adjust 
          Reply  Mon Jul 18 16:39:07 2022, Paco, Summary, LSC, FPMI locking procedure using REFL55 and AS55 DARM_07_18_2022_FMPI.pdfCARM_07_18_2022_FPMI.pdfMICH_07_18_2022_FPMI.pdffpmi_darm_nb_2022_07.pdf
             Reply  Mon Jul 18 21:41:42 2022, Anchal, Summary, LSC, FPMI locking procedure using REFL55 and AS55 
             Reply  Tue Aug 9 19:54:31 2022, yuta, Summary, LSC, FPMI locking tonight 
                Reply  Thu Aug 18 14:49:35 2022, Yehonathan, Summary, LSC, FPMI Sensitivity Sensitivity_Plot_1344133503.pdf
                   Reply  Thu Aug 18 18:10:49 2022, Koji, Summary, LSC, FPMI Sensitivity 
Message ID: 16968     Entry time: Fri Jul 1 08:50:48 2022     Reply to this: 17002
Author: yuta 
Type: Summary 
Category: LSC 
Subject: FPMI with REFL/AS55 trial 

[Anchal, Paco, Yuta]

We tried to lock FPMI with REFL55 and AS55 this week, but no success yet.
FPMI locks with POX11, POY11 and ASDC for MICH stably, but handing over to 55's couldn't be done yet.

What we did:
 - REFL55: Increased the whitening gain to 24dB. Demodulation phase tuned to minimize MICH signal in I when both arms are locked with POX and POY. REFL55 is noisier than AS55. Demodulation phase and amplitude of the signal seem to drift a lot also. Might need investigation.
 - AS55: Demodulation phase tuned to minimize MICH signal in I when both arms are locked with POX and POY. Whitening gain is 24dB.
 - Script for demodulation phase tuning lives in https://git.ligo.org/40m/scripts/-/blob/main/RFPD/getPhaseAngle.py
 - Locking MICH with REFL55 Q: Kicks BS much and not so stable probably because of noisy REFL55. Offtet also needs to be adjusted to lock MICH to dark fringe.
 - BS coil balancing: When MICH is "locked" with REFL55 Q, TRX drops rapidly and AS fringe gets worse, indicating BS coil balancing is not good. We balanced the coils by dithering POS with different coil output matrix gains to minimize oplev PIT and YAW output manually using LOCKINs.
 - Locking MICH with ASDC: Works nicely. Offset is set to -0.1 in MICH filter and reduced to -0.03 after lock acquisition.
 - ETMX/ETMY actuation balancing: We found that feedback signal to ETMX and ETMY at LSC output is unbalanced when locking with POX and POY. We dithered MC2 at 71 Hz, and checked feedback signals when Xarm/Yarm are locked to find out actuation efficiency imbalance. A gain of 2.9874 is put into C1:LSC-ETMX filter to balance ETMX/ETMY. I think we need to check this factor carefully again.
 - TRX and TRY: We normalized TRX and TRY to give 1 when arms are aligned. Before doing this, we also checked the alignment of TRX and TRY DC PDs (also reduced green scattering for TRY). Together with ETMX/ETMY balancing, this helped making filter gains the same for POX and POY lock to be 0.02 (See, also 40m/16888).
 - Single arm with REFL55/AS55: We checked that single arm locking with both REFL55_I and AS55_Q works. Single arm locking feeding back to MC2 also worked.
 - Handing over to REFL55/AS55: After locking Xarm and Yarm using POX to ETMX and POY to ETMY, MICH is locked with ASDC to BS. Handing over to REFL55_I for CARM using ETMX+ETMY and AS55_Q for DARM using -ETMX+ETMY was not successful. Changing an actuator for CARM to MC2 also didn't work. There might be an unstable point when turning off XARM/YARM filter modules and switching on DARM/CARM filter modules with a ramp time. We also need to re-investigate correct gains and signs for DARM and CARM. (Right now, gains are 0.02 for POX and POY, -0.02 for DARM with AS55_Q (-ETMX+ETMY), -0.02 for CARM with REFL55_I with MC2 are the best we found so far)
 
Next:
 - Measure ETMX and ETMY actuation efficiencies with Xarm/Yarm to balance the output matrix for DARM.
 - Measure optical gains of POX11, POY11, AS55 and REFL55 when FPMI is locked with POX/POY/ASDC to find out correct filter gains for them.
 - Make sure to measure OLTFs when doing above to correct for loop gains.
 - Lock CARM with POY11 to MC2, DARM with POX11 to ETMX. Use input matrix to hand over instead of changing filter modules from XARM/YARM to DARM/CARM.
 - Try using ALS to lock FPMI.

ELOG V3.1.3-